OpenStudy (anonymous):

How many L of O2 are needed to react completely with 63.0 L of H₂S at STP? (show your work) 2H₂S(g) + 3O₂(g) → 2SO₂(g) + 2H₂O(g)

OpenStudy (photon336):

STP = standard temperature and pressure, 293K 1 atm

OpenStudy (photon336):

\[pV =nRT \] we know R, T and V and P we just need n the number of moles of H2S \[\frac{ pV }{ RT } = n\] plug in our values \[\frac{ (1atm)(63.0L) }{ (273K)(0.08K ^{-1})Lmol ^{-1}atm } = 2.90 moles, H_{2}S\] one you know this number the number of moles which we just got. figuring out how many liters of O2 is easy. you can multiply by the molar ratio Do this and you'll get the number of moles of O2 \[2.90 mol H_{2}S *(\frac{ 3O_{2} }{ 2H_{2}S }) = moles of O_{2}\] then you can easily find the number of liters atstp

OpenStudy (anonymous):

can you help? how would i do that?

OpenStudy (photon336):

see above what I did try it out first the tell me what you would get